### v1

# LLM-Advisor: An LLM Benchmark for Cost-efficient Path Planning across Multiple Terrains

Abstract
Multi-terrain cost-efficient path planning is a crucial task in robot navigation, requiring the identification of a path from the start to the goal that not only avoids obstacles but also minimizes travel costs. This is especially crucial for real-world applications where robots need to navigate diverse terrains in outdoor environments, where recharging or refueling is difficult. However, there is very limited research on this topic. In this paper, we develop a prompt-based approach, LLM-Advisor, which leverages large language models (LLMs) as effective advisors for path planning. The LLM-Advisor selectively provides suggestions, demonstrating its ability to recognize when no modifications are necessary. When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm achieve greater cost efficiency. Since LLM-Advisor may occasionally lack common sense in their suggestions, we propose two hallucination-mitigation strategies. Furthermore, we experimentally verified that GPT-4o performs poorly in zero-shot path planning, even when terrain descriptions are clearly provided, demonstrating its low spatial awareness. We also experimentally demonstrate that using an LLM as an advisor is more effective than directly integrating it into the path-planning loop. Since LLMs may generate hallucinations, using LLMs in the loop of a search-based method (such as A*) may lead to a higher number of failed paths, demonstrating that our proposed LLM-Advisor is a better choice.
I Introduction
Path planning is a task of finding a path from a starting point to a goal while meeting specific conditions such as avoiding obstacles, minimizing travel distance or time, and adhering to other constraints [1]. This task is crucial in multiple fields such as robotics, autonomous navigation, industrial automation, and virtual environment navigation, as it directly impacts the efficiency, safety, and feasibility of operational systems [2]. Many path planning methods have been proposed in recent years, and they can effectively accomplish the planning task while ensuring the validity of the path in most cases [3, 4, 5, 6, 7, 8].
Existing methods for path planning are mainly categorized into search-based and sampling-based approaches. Typical search-based methods include Breadth-First Search (BFS) [9], Depth-First Search (DFS) [10], Best-First Search [11], Dijkstra’s algorithm [12], A* [13], and various A* variants [14, 15, 16, 17, 18, 19, 20]. On the other hand, typical sampling-based methods include Rapidly-exploring Random Tree (RRT) [21], RRT* [22], Informed RRT* [23], RRT-Connect [24], Dynamic RRT [25], Improved RRT [26], etc. However, these methods lack knowledge of the global environment, and they often generate paths with unnecessary detours, which may not be optimal.
Meanwhile, Large Language Models (LLMs) have a powerful capability for understanding global contexts [27, 28, 29, 30], which is especially beneficial in complex problem-solving domains such as path planning [31, 32, 33, 34]. These models have demonstrated the ability of LLMs to handle and reason over long-context inputs, providing global insights that reflect their understanding of the environment, such as identifying obstacles, agents, and the relative positions of goals. However, Ref. [35] found that LLMs perform poorly when dealing with complex, long-term planning and spatial reasoning tasks such as grid-based path planning. LLMs often generate invalid or impractical paths, leading to incomplete routes or collisions, indicating a capability gap when handling detailed spatial complexity. This paper aims to solve the above issues. We claim that by carefully designing the prompts, LLMs can serve as effective advisors for complex path-planning tasks.
In this paper, we first introduce a new task setting: cost-efficient path planning across multiple terrains, which is crucial for unmanned robots. While existing methods mainly consider the travel distance and time, the travel cost is also very important. Robots often operate in environments where recharging or refueling is challenging, such as remote areas, space, or disaster zones. Minimizing energy consumption allows robots to operate for longer periods without the need to recharge, thereby increasing their operational efficacy. Secondly, we propose an innovative path-planning method, LLM-Advisor. LLM-Advisor serves as a post-processing module for the path planning algorithms by leveraging global insights provided by large language models. It suggests improvements to the planned path when the optimal cost-efficient route has not been found. We experimentally found that: 1) By carefully designing the prompts, the proposed LLM-Advisor selectively provides suggestions only when necessary, demonstrating its ability to discern when adjustments are needed and when they are not, and 2) By integrating the proposed hallucination-mitigation strategies, the LLM-Advisor significantly improves its performance in suggesting paths. 3) When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm prove to be more cost-efficient. The contributions of this paper are threefold:
  • We propose a prompt-based method, LLM-Advisor, that enables LLMs to act as effective advisors after path planning is completed by other methods. Experimental results demonstrate the effectiveness of the proposed prompt and the validity of using LLMs as advisors for path planning.
  • We propose two hallucination mitigation strategies to further enhance the performance of LLM-Advisor.
  • Experimental results demonstrate that LLM-Advisor is capable of selectively making suggestions. Specifically, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm are more cost-efficient.
TABLE I: Terrain difficulty levels, traversal costs, and corresponding colors
Difficulty Level 0 1 2 3 4 5 6 7 8 9 10 11
Cost 1 3 3.5 4 4.5 5 5 6 6.5 7 8 $\infty$
Color
Figure 1: A visualization of samples from the created MultiTerraPath dataset
TABLE II: The definitions of traversal costs for the RUGD_v2 dataset
Color
Label Rock Grass Sign Water Pole Tree Bicycle Gravel Concrete Log Fence Rockbed
Traversal Cost $\infty$ 1.8 $\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 1.5 1 $\infty$ $\infty$ 3.0
Color
Label Bridge Dirt Table Mulch Sky Person Building Container (generic object) Alphalt Bush Sand Vehicle
Traversal Cost $\infty$ 2.2 $\infty$ 2.5 $\infty$ $\infty$ $\infty$ $\infty$ 1.2 $\infty$ 3.5 $\infty$
Figure 2: A visualization of samples from the created RUGD_v2 dataset
You are provided with a terrain cost grid description: {terrain_description} You are also provided with the real coordinates of start and goal points, and a description of path list computed by A*. Below is the data for each pair: **Start-End Pair **: - Start: {start} - End: {end} - A* path: {path_description} Analyzing the A* paths, evaluate whether the most cost-efficient path has already been found for each pair. IMPORTANT: **If not, you must suggest a path list, ensure the first point is the exact start coordinates and the last point is the exact end coordinates as provided, display every single coordinate pair**. IMPORTANT: **Make your reply strictly in the following format**: Requirements: If a path is optimal, the answer should be Yes. If a path is not optimal, the answer should be No, the suggested path coordinates (P). The list of coordinates (P) must start at the corresponding start coordinate and end at the corresponding end coordinate. The output format must strictly follow the format below, without any additional information. Output Format: Yes or No, No or [P] IMPORTANT: Only output the real list of coordinates for P. Do not include any extra explanations or text.
Figure 3: The designed prompt for LLM-Advisor
Refer to caption Terrain description: High-cost area with cost 5 is approximately located between grid coordinates (42, 110) and (354, 454). High-cost area with cost 4 is approximately located between grid coordinates (53, 268) and (131, 407). High-cost area with cost 6 is approximately located between grid coordinates (98, 113) and (350, 224). High-cost area with cost 3 is approximately located between grid coordinates (148, 38) and (264, 430). High-cost area with cost 8 is approximately located between grid coordinates (200, 346) and (271, 383). High-cost area with cost 7 is approximately located between grid coordinates (349, 102) and (413, 183). Obstacles are approximately located between grid coordinates (268, 270) and (312, 395). Path description (Detailed): Point 1 at (172, 88) has a terrain cost of 1. Point 2 at (173, 89) has a terrain cost of 1. Point 3 at (174, 90) has a terrain cost of 1. Point 4 at (175, 91) has a terrain cost of 1. Point 5 at (176, 92) has a terrain cost of 1. Point 6 at (177, 93) has a terrain cost of 1. Point 7 at (178, 94) has a terrain cost of 1. Point 8 at (179, 95) has a terrain cost of 1. … Point 285 at (437, 337) has a terrain cost of 5. Point 286 at (438, 337) has a terrain cost of 5. Point 287 at (439, 337) has a terrain cost of 5. Point 288 at (440, 337) has a terrain cost of 5. Point 289 at (441, 337) has a terrain cost of 5. Point 290 at (442, 337) has a terrain cost of 5. Point 291 at (443, 337) has a terrain cost of 5. The total cost of the path is 410.
Figure 4: Examples of terrain descriptions and detailed path descriptions for the A*-planned path in map 1 of the MultiTerraPath dataset
Retrieval-augmented strategy: You are provided with a terrain cost grid description based on historical tasks: {retrieved_terrain_description} Additionally, you have the start and goal coordinates along with computed A-star paths retrieved from similar tasks: {retrieved_path_description} Descriptive output: Requirements: If the path is optimal, respond with “Yes.” If the path is not optimal, respond with “No,” followed by a descriptive sentence indicating that the suggested path is not feasible, and then provide an alternative path. The description should clearly outline the movement from the start coordinate to the end coordinate, step by step. The output must strictly follow the format below, without any additional information. Output Format: Yes or No, No or from $(start_{x},start_{y})$ it goes to $(x_{1},y_{1})$, then $(x_{2},y_{2})$, …, finally, it arrives at $(end_{x},end_{y})$.
Figure 5: Hallucination-mitigation strategies
Figure 6: The A* algorithm search eight neighbors
Figure 7: Visualization of the A* searched paths
You are provided with a terrain cost grid description: {terrain_description} You are also provided with the real coordinates of start and goal points, **Start-End Pair**: - Start: {start} - End: {end} Suggest a list of intermediate target states, P, to help an A* algorithm plan the most cost-efficient path. The output format must strictly follow the format below, without any additional information. Output Format: [P] IMPORTANT: Only output the real list of coordinates for P. Do not include any extra explanations or text.
Figure 8: LLM-A* Prompt
II Method
In this section, we first introduce the dataset used in this paper. Then, we describe the details of the proposed LLM-Advisor.
II-A Datasets
MultiTerraPath Dataset. This paper introduces a synthetic dataset for evaluating model performance, MultiTerraPath. Our MultiTerraPath dataset contains 1,000 maps, each sized $500\times 500$. Each map includes several high-cost terrains and one or two impassable obstacles, with one start-to-end pair defined per image. The dataset can be used for tasks such as cost-efficient path planning and obstacle-avoidance path planning. The defined terrains, traversal costs, and their corresponding colors in the images are summarized in Table I. Fig. 1 presents some sample maps from the dataset. The coordinates of the start and end points of each pair are saved in the corresponding JSON file.
RUGD_v2 Dataset. The RUGD dataset [36] supports semantic understanding of unstructured outdoor scenes, including rough and challenging terrains. There are four scenes in the RUGD dataset: creek, park, trail, and village. The size of each image is $688\times 555$. Based on the RUGD dataset, we create one start-to-end pair per image, forming our RUGD_v2 dataset. The traversal costs for different terrains are synthetically defined by GPT-4o (Table II). Fig. 2 shows some examples from the dataset. Note that in real-world applications, it is crucial to conduct thorough experiments to accurately determine the traversal costs for each terrain type. These experiments help to fine-tune the cost values by considering factors such as robot types and the physical properties of the terrain.
II-B LLM-Advisor
The motivations for designing LLM-Advisor are twofold: 1) Existing path planning methods, whether search-based or sampling-based, focus on local regions when conducting path planning. This ensures a reasonable path but may result in unnecessary detours, which significantly affect the total traversal cost. 2) Although some LLM-powered path planning methods have already been developed [37, 34], Ref. [37] leveraged an LLM in an interactive manner, which requires several rounds conversations. Therefore, it is not suitable for real-time operations. Ref. [34] leveraged an LLM to enhance the efficiency of the A* algorithm by generating a candidate list during the search process. However, the LLM-A* approach [34] led to an increased rate of failures in finding valid paths. Considering the above two issues, we propose using LLMs as advisors as a viable solution. However, due to the LLM’s low spatial awareness and tendency for hallucinations, careful prompt design and the implementation of mitigation strategies are essential.
Main Prompt. Prompt design is a key component of the proposed LLM-Advisor. Specifically, when designing the main prompt, the terrain description is included as knowledge, as it provides critical information about the path planning environment. To address the issue of low spatial awareness, we also incorporate a detailed description of the planned path. By asking the LLMs to evaluate both the terrain and the path, we aim to enhance their spatial awareness and improve its performance. The code in Fig. 3 shows the designed prompt, where the GPT-4o is utilized. Examples of the terrain descriptions and detailed path descriptions used are presented in Fig. 4.
Hallucination-mitigation Strategies. Furthermore, we propose two mitigation strategies to alleviate the hallucinations of LLM-Advisor. Details are given in Fig. 5.
  • DescPath. Since LLM-Advisor relies heavily on the contextual understanding capabilities of LLMs, directly generating coordinate points of the path may introduce hallucinations. Therefore, we propose generating a descriptive path (DescPath) instead of just listing individual path coordinates.
  • RAG. We also propose a retrieval-augmented strategy (RAG) to mitigate these hallucinations. Specifically, for the MultiTerraPath dataset, we randomly retrieve one path planning example from the entire dataset. For the RUGD_v2 dataset, we randomly select a path-planning example from the same scene and integrate the retrieved information into the original prompt.
III Experimental Results
III-A Experimental Setup and Evaluation Metrics
We utilize a GPT-4o to implement our proposed prompt method. We applied the proposed LLM-Advisor to check the planned path generated by A* [13], RRT* [22], and LLM-A* [34]. A* is a search-based method, RRT* is a sampling-based method, and LLM-A* is an LLM-based method. Note that the experiments with LLM-A* [34] also use a GPT-4o. To further evaluate the effectiveness of the LLM-Advisor, we also compare it with LLM-A* [34] and GPT-4o.
Overview of A*. A* is a search-based path planning algorithm. It evaluates the eight neighboring points around the current position and determines the most cost-efficient path by selecting the next step with the lowest combined cost of traversal and estimated remaining distance(Fig. 6). Therefore, A* generates reasonable paths, as illustrated in Fig. 7. However, the generated path often includes unnecessary detours. Although A* avoids obstacles, the overall path may not be the most cost-efficient.
TABLE III: Evaluation of planned paths across different methods on the MultiTerraPath dataset using the proposed LLM-Advisor
Method Terrain description Path description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
Detailed Brief
LLM-Advisor + A* [13] - 1,000 1,000 612 432 70.59
LLM-Advisor + RRT* [22] - 1,000 1,000 737 512 69.47
LLM-Advisor + LLM-A* [34] - 1,000 1,000 770 606 78.70
TABLE IV: Evaluation of planned paths across different methods on the RUGD_v2 dataset using the proposed LLM-Advisor
Method Scenes $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
LLM-Advisor + A* Creek 1000 1000 876 580 66.21
Park 1000 1000 860 720 83.72
Trail 1000 1000 620 540 87.10
Village 1000 1000 780 560 71.79
TABLE V: Comparison of our LLM-Advisor with other LLM-based methods
Method Terrain description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Successful}}$ $N_{\text{Improved}}$ $N_{\text{Equal}}$ $N_{\text{Deteriorated}}$ $IR$ (%) $\uparrow$ FPS $\uparrow$
A* [13] - 1,000 1,000 1,000 - - - - 9.39
GPT-4o 1,000 1,000 1,000 208 0 792 20.80 1.21
LLM-A* [34] 1,000 1,000 938 152 538 248 69.00 3.10
LLM-Advisor + A* 1,000 1,000 1,000 432 393 175 82.50 5.86
Figure 9: Visualization of the planned paths by GPT-4o, A*, LLM-A*, and LLM-Advisor + A*. Black bold numbers indicate the favorable suggestions, while red bold numbers highlight less favorable suggestions. All other numbers represent cases where no suggestions were made. Lower costs indicate better performance
Figure 10: Visualization of experimental results on RUGD_v2 dataset
TABLE VI: Ablation studies
Method Terrain description Path description RAG Descriptive output $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
Detailed Brief
LLM-Advisor + A* - - - 1,000 1,000 602 130 21.59
- - - 1,000 1,000 607 352 57.99
- - 1,000 1,000 610 413 67.70
- - 1,000 1,000 609 399 65.52
- 1,000 1,000 612 432 70.59
Path description (Brief): From (88, 172) to (305, 370) the terrain cost is 1. From (306, 371) to (306, 371) the terrain cost is 9223372036854775807. From (307, 372) to (337, 413) the terrain cost is 1. From (337, 414) to (337, 443) the terrain cost is 5. The total cost of the path is -9223372036854775399.
Figure 11: Path description (Brief)
Figure 12: Visualization of experimental results after integrating the proposed hallucination-mitigation strategies
Overview of RRT*. RRT* is a sampling-based path planning algorithm that extends RRT [21]. It randomly samples points and connects each new point to the nearest neighbor in the tree. Unlike RRT, RRT* evaluates the cost from all nearby nodes and selects the lowest-cost connection. After connecting, it rewires the tree whenever a more efficient path is found. Over time, this iterative rewiring process improves the overall path quality. Although time-consuming, RRT* generates reasonable paths, which is why we use our LLM-Advisor to evaluate its results, aiming to assess its effectiveness as a validation tool for precomputed reasonable paths.
Overview of LLM-A*. LLM-A* leverages LLMs to provide a candidate list for the A* algorithm. The main idea of LLM-A* is to improve search efficiency in obstacle-avoidance path planning. By restricting the search to the candidate list provided by LLM, LLM-A* can significantly decrease the search time. The prompt used for experiments using LLM-A* [34] is given in Fig. 8.
Since LLM-Advisor functions as an advisor, we assess its performance using two metrics: Relative Precision ($RP$) and Improvement Ratio ($IR$). The definition of $RP$ is provided in Eq. 1, and the definition of $IR$ is given in Eq. 2. $RP$ represents the proportion of improved paths within the suggestions made by LLM-Advisor, while $IR$ indicates the percentage of paths that are equal to or better than the original ones after incorporating LLM-Advisor. $IR$ can reveal how many paths are valid and effective.
$RP=\frac{N_{\text{Improved}}}{N_{\text{Suggested}}},$ (1)
where $N_{\text{Improved}}$ denotes the number of improved paths and $N_{\text{Suggested}}$ denotes the number of suggested paths.
$IR=\frac{N_{\text{Improved}}+N_{\text{Equal}}}{N_{\text{Path}}},$ (2)
where $N_{\text{Equal}}$ denotes the number of paths that are identical to the A*-planned path, and $N_{\text{Path}}$ represents the total number of planned paths.
III-B Main Results
Main Comparisons. Table III presents the main comparison results for the MultiTerraPath dataset. Most suggestions made by the proposed LLM-Advisor are more cost-efficient. It achieves an $RP$ of 70.59% when making suggestions for the A*-planned path, 69.47% for the RRT*-planned path, and 78.70% for the LLM-A*-planned path. This demonstrates that, with detailed path descriptions, LLM-Advisor can effectively serve as a reliable validation tool for planned paths generated by other methods. Moreover, the proposed LLM-Advisor is versatile. It can improve the previous approaches regardless of whether they are search-based, sampling-based, or LLM-based. Table IV shows experimental results for the the RUGD_v2 dataset.
Table V presents a comparison of LLM-Advisor+A* with A*, LLM-A* [34] and GPT-4o. The best performance for each evaluation metric is bolded, while the second-best is underlined. A* successfully generated valid paths for all cases. Since the proposed LLM-Advisor functions as an advisor, evaluating the paths generated by A*, it also successfully validated all paths. GPT-4o successfully planned all paths, although most were suboptimal or invalid. However, it was still able to generate a path. In LLM-A [34], GPT-4o suggests a target list, providing a sequence of intermediate target states for A* to consider during path planning. If the generated target list contains points within the obstacle region, the neighbor search process cannot proceed. As shown in Table V, there were 62 paths that were not successfully planned using LLM-A*. When considering the Improvement Ratio (IR), LLM-Advisor+A* achieved a higher IR, indicating better performance. The GPT-4o has the lowest IR, indicating that GPT-4o does not have ability to do zero-shot path planning. Moreover, since our LLM-Advisor acts as a postprocessor, it is possible to involve humans in the verification process, allowing us to decide whether to accept the paths suggested by the LLM-Advisor. Incorporating human judgment in this process can further improve the quality of the planned paths.
To evaluate the computational efficiency of different path planning methods, we report frames per second (FPS) for each approach in Table V. FPS quantifies the number of paths processed per second, where a higher value indicates greater efficiency. All methods are executed entirely on the CPU. Among them, LLM-Advisor + A* selectively makes suggestions only when an improvement is needed, minimizing unnecessary LLM computations and achieving a higher FPS compared to other LLM-based methods.
Visualizations. Fig. 9 presents some path planning results. As shown in the figure, our LLM-Advisor successfully suggests more cost-efficient paths. Fig. 10 provides visualization results for the RUGD_v2 dataset.
Ablation Studies. Table VI shows some ablation study results. We evaluate LLM-Advisor with both brief and detailed path descriptions to demonstrate the effectiveness of the proposed prompt. One example of brief path description is given in Fig. 11. As shown, the level of detail in path descriptions significantly impacts final performance. By using detailed path descriptions, the LLM-Advisor can more accurately evaluate the A*-planned path and make a higher-quality suggestion. We also found that it has only a limited influence on the number of suggestions. This indicates that the ability of LLM-Advisor to suggest improvements to planned paths is not affected much by the path descriptions. Moreover, using the proposed hallucination-mitigation strategies, the LLM-Advisor achieves higher RP. Fig. 12 shows the visualization results.
IV Future Work
The proposed LLM-Advisor operates as a zero-shot model, relying heavily on the performance of general-purpose pretrained LLMs. Without task-specific training, it leverages the inherent knowledge and reasoning capabilities of these models to provide guidance and insights. However, its performance may be limited when dealing with domain-specific challenges. As a next step, we plan to fine-tune the model using task-relevant data to enhance its accuracy, adaptability, and effectiveness in path planning and related tasks.
V Conclusions
This paper introduced a novel prompt-based approach, LLM-Advisor, which achieves high relative precision and improvement ratio in evaluating paths planned by existing methods. The LLM-Advisor is versatile and can be used to verify paths generated by any planing method. It provides suggestions when existing methods fail to identify the most cost-efficient path. Furthermore, we propose two strategies to mitigate hallucinations. Our experiments on created MultiTerraPath and RUGD_v2 datasets demonstrate the effectiveness of LLM-Advisor. We also experimentally found that GPT-4o struggles with effective path planning based solely on terrain descriptions. We believe that our work offers valuable insights for the path planning community, particularly for researchers investigating the integration of LLMs into path planning.
Acknowledgments
This paper is partially financially supported by JSPS KAKENHI Grant Number 24K20787, Fund to Support Female Researchers with Their Research Skills Training of The University of Tokyo, and Institute for AI and Beyond of The University of Tokyo.

### v2

@@
 Abstract
-Multi-terrain cost-efficient path planning is a crucial task in robot navigation, requiring the identification of a path from the start to the goal that not only avoids obstacles but also minimizes travel costs.
-This is especially crucial for real-world applications where robots need to navigate diverse terrains in outdoor environments, where recharging or refueling is difficult.
-However, there is very limited research on this topic.
-In this paper, we develop a prompt-based approach, LLM-Advisor, which leverages large language models (LLMs) as effective advisors for path planning.
-The LLM-Advisor selectively provides suggestions, demonstrating its ability to recognize when no modifications are necessary.
-When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm achieve greater cost efficiency.
-Since LLM-Advisor may occasionally lack common sense in their suggestions, we propose two hallucination-mitigation strategies.
-Furthermore, we experimentally verified that GPT-4o performs poorly in zero-shot path planning, even when terrain descriptions are clearly provided, demonstrating its low spatial awareness.
-We also experimentally demonstrate that using an LLM as an advisor is more effective than directly integrating it into the path-planning loop.
-Since LLMs may generate hallucinations, using LLMs in the loop of a search-based method (such as A*) may lead to a higher number of failed paths, demonstrating that our proposed LLM-Advisor is a better choice.
+Cost-efficient path planning across multiple terrains is a crucial task in robot navigation, requiring the identification of a path from the start to the goal that not only avoids obstacles but also minimizes the overall travel cost.
+This is especially crucial for real-world applications where robots need to navigate diverse terrains in outdoor environments with limited opportunities for recharging or refueling.
+Despite its practical importance, cost-efficient path planning across heterogeneous terrains has received relatively limited attention in prior work.
+In this paper, we propose LLM-Advisor, a prompt-based, planner-agnostic framework that leverages large language models (LLMs) as non-decisive post-processing advisors for cost refinement, without modifying the underlying planner.
+While we observe that LLMs may occasionally produce implausible suggestions, we introduce two effective hallucination-mitigation strategies.
+We further introduce two datasets, MultiTerraPath and RUGD_v2, for systematic evaluation of cost-efficient path planning.
+Extensive experiments reveal that state-of-the-art LLMs, including GPT-4o, GPT-4-turbo, Gemini-2.5-Flash, and Claude-Opus-4, perform poorly in zero-shot terrain-aware path planning, highlighting their limited spatial reasoning capability.
+In contrast, the proposed LLM-Advisor (with GPT-4o) improves cost efficiency for 72.37% of A*-planned paths, 69.47% of RRT*-planned paths, and 78.70% of LLM-A*-planned paths.
+On the MultiTerraPath dataset, LLM-Advisor demonstrates stronger performance on the hard subset, further validating its applicability to real-world scenarios.
 I Introduction
-Path planning is a task of finding a path from a starting point to a goal while meeting specific conditions such as avoiding obstacles, minimizing travel distance or time, and adhering to other constraints [1].
-This task is crucial in multiple fields such as robotics, autonomous navigation, industrial automation, and virtual environment navigation, as it directly impacts the efficiency, safety, and feasibility of operational systems [2].
-Many path planning methods have been proposed in recent years, and they can effectively accomplish the planning task while ensuring the validity of the path in most cases [3, 4, 5, 6, 7, 8].
-Existing methods for path planning are mainly categorized into search-based and sampling-based approaches.
-Typical search-based methods include Breadth-First Search (BFS) [9], Depth-First Search (DFS) [10], Best-First Search [11], Dijkstra’s algorithm [12], A* [13], and various A* variants [14, 15, 16, 17, 18, 19, 20].
+Path planning is a task of finding a path from a starting point to a goal while meeting specific conditions such as avoiding obstacles, minimizing travel distance or time, and meeting other constraints [31, 23, 24, 11, 28].
+This task is crucial in multiple fields such as robotics, autonomous navigation, industrial automation, and virtual environment navigation, as it directly impacts the operational efficiency, safety, and feasibility of systems [6].
+Many path planning methods have been proposed in recent years, and they can effectively accomplish the planning task while ensuring the validity of the path in most cases [13, 3, 18, 12, 5].
+Existing methods are mainly search-based and sampling-based approaches.
+Typical search-based methods include Breadth-First Search (BFS) [30], Depth-First Search (DFS) [15], Best-First Search [19], Dijkstra’s algorithm [14], A* [4], and various A* variants [17, 8].
 On the other hand, typical sampling-based methods include Rapidly-exploring Random Tree (RRT) [10], RRT* [7], Informed RRT* [2], RRT-Connect [9], Dynamic RRT [29], Improved RRT [22], etc.
-However, these methods lack knowledge of the global environment, and they often generate paths with unnecessary detours, which may not be optimal.
-Meanwhile, Large Language Models (LLMs) have a powerful capability for understanding global contexts [27, 28, 29, 30], which is especially beneficial in complex problem-solving domains such as path planning [31, 32, 33, 34].
+Despite their effectiveness, these planners perform global optimization over discretized graph representations, whose structure and connectivity are constrained by map resolution, neighborhood definitions, and hand-crafted cost models.
+For search-based planners, heuristic-guided expansions operate on predefined graph neighborhoods, which limits the expressiveness of the search space and its ability to capture large-scale structural patterns or long-range cost trends present in the continuous environment.
+As a result, even with access to the full cost map, the resulting paths may still exhibit unnecessary detours or suboptimal cost trade-offs.
+Meanwhile, Large Language Models (LLMs) have strong capabilities for understanding global contexts [27], which is particularly beneficial in complex problem-solving domains such as path planning [21, 20, 16].
 These models have demonstrated the ability of LLMs to handle and reason over long-context inputs, providing global insights that reflect their understanding of the environment, such as identifying obstacles, agents, and the relative positions of goals.
-However, Ref.
+However, Aghzal et al.
 [1] found that LLMs perform poorly when dealing with complex, long-term planning and spatial reasoning tasks such as grid-based path planning.
 LLMs often generate invalid or impractical paths, leading to incomplete routes or collisions, indicating a capability gap when handling detailed spatial complexity.
-This paper aims to solve the above issues.
+This paper aims to address these issues.
 We claim that by carefully designing the prompts, LLMs can serve as effective advisors for complex path-planning tasks.
-In this paper, we first introduce a new task setting: cost-efficient path planning across multiple terrains, which is crucial for unmanned robots.
-While existing methods mainly consider the travel distance and time, the travel cost is also very important.
-Robots often operate in environments where recharging or refueling is challenging, such as remote areas, space, or disaster zones.
-Minimizing energy consumption allows robots to operate for longer periods without the need to recharge, thereby increasing their operational efficacy.
-Secondly, we propose an innovative path-planning method, LLM-Advisor.
-LLM-Advisor serves as a post-processing module for the path planning algorithms by leveraging global insights provided by large language models.
-It suggests improvements to the planned path when the optimal cost-efficient route has not been found.
-We experimentally found that: 1) By carefully designing the prompts, the proposed LLM-Advisor selectively provides suggestions only when necessary, demonstrating its ability to discern when adjustments are needed and when they are not, and 2) By integrating the proposed hallucination-mitigation strategies, the LLM-Advisor significantly improves its performance in suggesting paths.
-3) When suggestions are made, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm prove to be more cost-efficient.
-The contributions of this paper are threefold:
+In this paper, we introduce a new task: cost-efficient path planning across multiple terrains, a critical challenge for unmanned robots.
+Unlike existing methods that mainly optimize distance or time, we emphasize travel cost.
+Given a path represented as a sequence of connected segments, the total travel cost is defined as the sum over all path segments, where each segment cost is computed as the product of its Euclidean length and the corresponding terrain cost.
+This formulation is particularly relevant for robots operating in environments where recharging is difficult (e.g., remote areas, space, and disaster zones).
+Reducing energy use directly extends operational time and effectiveness.
+However, achieving cost-efficient navigation remains challenging in practice.
+Although classical planners such as A* are globally optimal with respect to a given discrete graph formulation, in practical robot navigation settings they typically operate on discretized maps with limited connectivity, due to sensing resolution, computational constraints, or map availability.
+As a result of discretization effects and limited connectivity, the resulting paths may still be suboptimal with respect to the underlying continuous terrain and long-range cost structure.
+This motivates the need for an additional mechanism that can reason over the global terrain layout and propose path refinements that are not representable within the planner’s original discrete search space, while leaving A*’s optimality guarantees on the discrete graph intact.
+Specifically, in our experimental setting, A* operates on a discretized grid graph with fixed 8-connected connectivity and uses the Manhattan distance to the goal as the heuristic function.
+We then propose LLM-Advisor, a post-processing module that leverages large language models to refine paths generated by classical planners.
+It selectively suggests improvements only when needed and incorporates hallucination-mitigation strategies, achieving significant gains: 72.37% improved for A*, 69.47% for RRT*, and 78.70% for LLM-A*.
+Unlike prior LLM approaches that rely on end-to-end generation or interactive queries, our method uses LLMs in an advisory role.
+With careful prompt design and hallucination control, LLM-Advisor improves cost efficiency without retraining or modifying the planner.
+Importantly, we do not aim to outperform optimally designed continuous or high-connectivity planners; instead, we target practical scenarios with fixed discretization, limited connectivity, or legacy implementations, where LLM-Advisor provides a planner-agnostic and low-risk mechanism for cost refinement.
+In summary, the contributions of this paper are threefold:
   • We propose a prompt-based method, LLM-Advisor, that enables LLMs to act as effective advisors after path planning is completed by other methods.
-Experimental results demonstrate the effectiveness of the proposed prompt and the validity of using LLMs as advisors for path planning.
   • We propose two hallucination mitigation strategies to further enhance the performance of LLM-Advisor.
   • Experimental results demonstrate that LLM-Advisor is capable of selectively making suggestions.
-Specifically, 70.59% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm are more cost-efficient.
-TABLE I: Terrain difficulty levels, traversal costs, and corresponding colors
+Specifically, 72.37% of the paths suggested for the A* algorithm, 69.47% for the RRT* algorithm, and 78.70% for the LLM-A* algorithm are more cost-efficient.
+The proposed hallucination-mitigation strategies significantly reduce hallucinations.
+TABLE I: Terrain difficulty levels, travel costs, and corresponding colors.
 Difficulty Level 0 1 2 3 4 5 6 7 8 9 10 11
@@
 Color
-Figure 1: A visualization of samples from the created MultiTerraPath dataset
-TABLE II: The definitions of traversal costs for the RUGD_v2 dataset
+Figure 1: MultiTerraPath samples.
+Easy maps contain separated convex high-cost regions, while hard maps include irregular, connected terrains with no purely zero-cost paths, requiring cost-aware trade-offs.
+Figure 2: A visualization of samples from the created RUGD_v2 dataset.
+TABLE II: The definitions of travel costs for the RUGD_v2 dataset.
 Color
 Label Rock Grass Sign Water Pole Tree Bicycle Gravel Concrete Log Fence Rockbed
-Traversal Cost $\infty$ 1.8 $\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 1.5 1 $\infty$ $\infty$ 3.0
+Travel Cost $\infty$ 1.8 $\infty$ $\infty$ $\infty$ $\infty$ $\infty$ 1.5 1 $\infty$ $\infty$ 3.0
 Color
 Label Bridge Dirt Table Mulch Sky Person Building Container (generic object) Alphalt Bush Sand Vehicle
-Traversal Cost $\infty$ 2.2 $\infty$ 2.5 $\infty$ $\infty$ $\infty$ $\infty$ 1.2 $\infty$ 3.5 $\infty$
-Figure 2: A visualization of samples from the created RUGD_v2 dataset
-You are provided with a terrain cost grid description: {terrain_description} You are also provided with the real coordinates of start and goal points, and a description of path list computed by A*.
-Below is the data for each pair: **Start-End Pair **: - Start: {start} - End: {end} - A* path: {path_description} Analyzing the A* paths, evaluate whether the most cost-efficient path has already been found for each pair.
+Travel Cost $\infty$ 2.2 $\infty$ 2.5 $\infty$ $\infty$ $\infty$ $\infty$ 1.2 $\infty$ 3.5 $\infty$
+Figure 3: The pipeline of proposed LLM-Advisor.
+You are provided with a terrain cost grid description: {terrain_description}.
+You are also provided with the real coordinates of the start and goal points, and a description of the path list computed by {existing path planning method}.
+Below is the data for each pair: **Start-End Pair **: - Start: {start} - End: {end} - {existing path planning method}-planned path: {path_description} Analyzing the {existing path planning method}-planned path, evaluate whether the most cost-efficient path has already been found for each pair.
 IMPORTANT: **If not, you must suggest a path list, ensure the first point is the exact start coordinates and the last point is the exact end coordinates as provided, display every single coordinate pair**.
@@
 Do not include any extra explanations or text.
-Figure 3: The designed prompt for LLM-Advisor
+Figure 4: The designed prompt for LLM-Advisor.
 Refer to caption Terrain description: High-cost area with cost 5 is approximately located between grid coordinates (42, 110) and (354, 454).
-High-cost area with cost 4 is approximately located between grid coordinates (53, 268) and (131, 407).
-High-cost area with cost 6 is approximately located between grid coordinates (98, 113) and (350, 224).
-High-cost area with cost 3 is approximately located between grid coordinates (148, 38) and (264, 430).
-High-cost area with cost 8 is approximately located between grid coordinates (200, 346) and (271, 383).
-High-cost area with cost 7 is approximately located between grid coordinates (349, 102) and (413, 183).
-Obstacles are approximately located between grid coordinates (268, 270) and (312, 395).
+High-cost area with cost 4 is approximately located between grid coordinates (53, 268) and (131, 407)… Obstacles are approximately located between grid coordinates (268, 270) and (312, 395).
 Path description (Detailed): Point 1 at (172, 88) has a terrain cost of 1.
 Point 2 at (173, 89) has a terrain cost of 1.
-Point 3 at (174, 90) has a terrain cost of 1.
-Point 4 at (175, 91) has a terrain cost of 1.
-Point 5 at (176, 92) has a terrain cost of 1.
-Point 6 at (177, 93) has a terrain cost of 1.
-Point 7 at (178, 94) has a terrain cost of 1.
-Point 8 at (179, 95) has a terrain cost of 1.
-… Point 285 at (437, 337) has a terrain cost of 5.
-Point 286 at (438, 337) has a terrain cost of 5.
-Point 287 at (439, 337) has a terrain cost of 5.
-Point 288 at (440, 337) has a terrain cost of 5.
-Point 289 at (441, 337) has a terrain cost of 5.
-Point 290 at (442, 337) has a terrain cost of 5.
+Point 3 at (174, 90) has a terrain cost of 1 … Point 290 at (442, 337) has a terrain cost of 5.
 Point 291 at (443, 337) has a terrain cost of 5.
 The total cost of the path is 410.
-Figure 4: Examples of terrain descriptions and detailed path descriptions for the A*-planned path in map 1 of the MultiTerraPath dataset
-Retrieval-augmented strategy: You are provided with a terrain cost grid description based on historical tasks: {retrieved_terrain_description} Additionally, you have the start and goal coordinates along with computed A-star paths retrieved from similar tasks: {retrieved_path_description} Descriptive output: Requirements: If the path is optimal, respond with “Yes.” If the path is not optimal, respond with “No,” followed by a descriptive sentence indicating that the suggested path is not feasible, and then provide an alternative path.
+Figure 5: Examples of terrain descriptions and detailed path descriptions for the A*-planned path in map 1 of the MultiTerraPath dataset.
+The terrain descriptions are automatically generated by grouping grid cells with similar cost values, extracting their bounding boxes, and converting them into concise text summaries.
+This process is fully automatic and scales linearly with map size.
+Retrieval-augmented strategy: You are provided with a terrain cost grid description based on historical tasks: {retrieved_terrain_description}.
+Additionally, you have the start and goal coordinates along with the computed {existing path planning method}-planned path retrieved from similar tasks: {retrieved_path_description}.
+Descriptive output: Requirements: If the path is optimal, respond with “Yes.” If the path is not optimal, respond with “No,” followed by a descriptive sentence indicating that the suggested path is not feasible, and then provide an alternative path.
 The description should clearly outline the movement from the start coordinate to the end coordinate, step by step.
@@
 Output Format: Yes or No, No or from $(start_{x},start_{y})$ it goes to $(x_{1},y_{1})$, then $(x_{2},y_{2})$, …, finally, it arrives at $(end_{x},end_{y})$.
-Figure 5: Hallucination-mitigation strategies
-Figure 6: The A* algorithm search eight neighbors
-Figure 7: Visualization of the A* searched paths
+Figure 6: Hallucination-mitigation strategies.
 You are provided with a terrain cost grid description: {terrain_description} You are also provided with the real coordinates of start and goal points, **Start-End Pair**: - Start: {start} - End: {end} Suggest a list of intermediate target states, P, to help an A* algorithm plan the most cost-efficient path.
@@
 Do not include any extra explanations or text.
-Figure 8: LLM-A* Prompt
+Figure 7: LLM-A* prompt.
 II Method
-In this section, we first introduce the dataset used in this paper. Then, we describe the details of the proposed LLM-Advisor.
 II-A Datasets
 MultiTerraPath Dataset.
-This paper introduces a synthetic dataset for evaluating model performance, MultiTerraPath.
-Our MultiTerraPath dataset contains 1,000 maps, each sized $500\times 500$.
-Each map includes several high-cost terrains and one or two impassable obstacles, with one start-to-end pair defined per image.
+This paper introduces MultiTerraPath, a synthetic dataset designed for evaluating model performance in cost-aware path planning.
+MultiTerraPath contains 2,000 maps, each with a resolution of $500\times 500$, comprising an easy subset with 1,000 maps and a hard subset with 1,000 maps.
+Easy map contains convex high-cost regions that are well separated by zero-cost areas, allowing straightforward low-cost routing.
+Hard map, in contrast, features irregular and concave terrain regions (e.g., corridor- or ribbon-shaped structures), allow terrain regions to be connected or overlapping to better mimic real-world continuous landscapes, and include cases where no purely zero-cost path exists, thereby forcing explicit cost-aware trade-offs.
 The dataset can be used for tasks such as cost-efficient path planning and obstacle-avoidance path planning.
-The defined terrains, traversal costs, and their corresponding colors in the images are summarized in Table I.
+The defined terrains, travel costs, and their corresponding colors in the images are summarized in Table I.
 Fig.
@@
 RUGD_v2 Dataset.
-The RUGD dataset [36] supports semantic understanding of unstructured outdoor scenes, including rough and challenging terrains.
-There are four scenes in the RUGD dataset: creek, park, trail, and village.
-The size of each image is $688\times 555$.
-Based on the RUGD dataset, we create one start-to-end pair per image, forming our RUGD_v2 dataset.
-The traversal costs for different terrains are synthetically defined by GPT-4o (Table II).
+The RUGD dataset [25] provides pixel-level semantic annotations for unstructured outdoor environments, covering four scenes: creek, park, trail, and village, with image resolution $688\times 555$.
+We create one start–goal pair per image to construct our RUGD_v2 dataset.
+Terrain-dependent travel costs are synthetically assigned using GPT-4o (Table II), and regions with infinite cost are treated as impassable obstacles.
 Fig.
-2 shows some examples from the dataset.
-Note that in real-world applications, it is crucial to conduct thorough experiments to accurately determine the traversal costs for each terrain type.
-These experiments help to fine-tune the cost values by considering factors such as robot types and the physical properties of the terrain.
+2 shows sample images.
+In real applications, terrain costs should be calibrated through physical experiments considering robot type and terrain properties.
 II-B LLM-Advisor
-The motivations for designing LLM-Advisor are twofold: 1) Existing path planning methods, whether search-based or sampling-based, focus on local regions when conducting path planning.
-This ensures a reasonable path but may result in unnecessary detours, which significantly affect the total traversal cost.
+The motivations for designing LLM-Advisor are twofold: 1) Existing path planning methods, whether search-based or sampling-based, operate on discretized maps with limited connectivity, due to sensing resolution, computational constraints, or map availability.
+While this often yields feasible paths, it can lead to unnecessary detours, increasing the total travel cost.
 2) Although some LLM-powered path planning methods have already been developed [26, 16], Ref.
-[37] leveraged an LLM in an interactive manner, which requires several rounds conversations.
+[26] leveraged an LLM in an interactive manner, which requires multiple rounds of conversation.
 Therefore, it is not suitable for real-time operations.
@@
 [16] leveraged an LLM to enhance the efficiency of the A* algorithm by generating a candidate list during the search process.
-However, the LLM-A* approach [34] led to an increased rate of failures in finding valid paths.
-Considering the above two issues, we propose using LLMs as advisors as a viable solution.
-However, due to the LLM’s low spatial awareness and tendency for hallucinations, careful prompt design and the implementation of mitigation strategies are essential.
+However, the LLM-A* approach [16] exhibited a higher failure rate in finding valid paths.
+To address these two issues, we propose using LLMs as non-decisive advisors.
+However, due to the limited spatial awareness and tendency for hallucinations in LLMs, careful prompt design and the implementation of mitigation strategies are essential.
+The overall pipeline of LLM-Advisor is provided in Fig.
+3.
 Main Prompt.
 Prompt design is a key component of the proposed LLM-Advisor.
-Specifically, when designing the main prompt, the terrain description is included as knowledge, as it provides critical information about the path planning environment.
-To address the issue of low spatial awareness, we also incorporate a detailed description of the planned path.
-By asking the LLMs to evaluate both the terrain and the path, we aim to enhance their spatial awareness and improve its performance.
+Specifically, the main prompt incorporates terrain descriptions as prior knowledge, providing critical information about the path planning environment.
+To address the issue of limited spatial awareness, we also incorporate a detailed description of the planned path.
+By prompting the LLMs to evaluate both the terrain and the path, we aim to enhance their spatial awareness and overall performance.
 The code in Fig.
@@
 Examples of the terrain descriptions and detailed path descriptions used are presented in Fig.
-4.
-Hallucination-mitigation Strategies. Furthermore, we propose two mitigation strategies to alleviate the hallucinations of LLM-Advisor. Details are given in Fig. 5.
+5.
+To ensure generalizability, our prompt design follows a modular template that separates terrain encoding, path description, and evaluation directive.
+This structure enables easy adaptation across tasks and environments.
+Moreover, the same prompt format was used successfully on both the synthetic MultiTerraPath and real-scene RUGD v2 datasets, suggesting that the prompt structure generalizes well without manual tuning per environment.
+Hallucination-mitigation Strategies. Furthermore, we propose two strategies to alleviate the hallucinations in LLM-Advisor. Details are given in Fig. 6.
   • DescPath.
-Since LLM-Advisor relies heavily on the contextual understanding capabilities of LLMs, directly generating coordinate points of the path may introduce hallucinations.
+Since LLM-Advisor relies heavily on the contextual reasoning capabilities of LLMs, directly generating coordinate points of the path may introduce hallucinations.
 Therefore, we propose generating a descriptive path (DescPath) instead of just listing individual path coordinates.
@@
 III-A Experimental Setup and Evaluation Metrics
-We utilize a GPT-4o to implement our proposed prompt method.
 We applied the proposed LLM-Advisor to check the planned path generated by A* [4], RRT* [7], and LLM-A* [16].
 A* is a search-based method, RRT* is a sampling-based method, and LLM-A* is an LLM-based method.
-Note that the experiments with LLM-A* [34] also use a GPT-4o.
-To further evaluate the effectiveness of the LLM-Advisor, we also compare it with LLM-A* [34] and GPT-4o.
-Overview of A*.
-A* is a search-based path planning algorithm.
-It evaluates the eight neighboring points around the current position and determines the most cost-efficient path by selecting the next step with the lowest combined cost of traversal and estimated remaining distance(Fig.
-6).
-Therefore, A* generates reasonable paths, as illustrated in Fig.
+The prompt used for LLM-A* [16] is given in Fig.
 7.
-However, the generated path often includes unnecessary detours.
-Although A* avoids obstacles, the overall path may not be the most cost-efficient.
-TABLE III: Evaluation of planned paths across different methods on the MultiTerraPath dataset using the proposed LLM-Advisor
-Method Terrain description Path description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
+To further evaluate the effectiveness of LLM-Advisor, we also compare it with LLM-A* [16], GPT-4o, GPT-4-turbo, Gemini-2.5-flash, and Claude-Opus-4.
+We adopt the same cost calculation method as A* [4] for all compared approaches.
+TABLE III: Evaluation of planned paths across different methods on the MultiTerraPath dataset using the proposed LLM-Advisor.
+Method Terrain description Path description Easy subset Hard subset
+Detailed Brief RP (%) $\uparrow$ SR (%) $\uparrow$ SCP (%) $\uparrow$ RP (%) $\uparrow$ SR (%) $\uparrow$ SCP (%) $\uparrow$
+LLM-Advisor (GPT-4o) + A* [4] - 72.37 99.10 94.10 74.56 99.30 94.32
+LLM-Advisor (GPT-4o) + RRT* [7] - 69.47 96.00 88.25 73.21 96.70 91.58
+LLM-Advisor (GPT-4o) + LLM-A* [16] - 78.70 97.00 84.47 81.39 97.70 89.91
+TABLE IV: Evaluations on the RUGD_v2 dataset using the proposed LLM-Advisor. $N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$.
+Method Scenes $N_{\text{Suggested}}$ $N_{\text{Improved}}$ RP (%) $\uparrow$ SR(%) $\uparrow$ SCP(%) $\uparrow$
+LLM-Advisor (GPT-4o) + A* Creek 876 580 66.21 94.70 84.39
+Park 860 720 83.72 99.10 89.11
+Trail 620 540 87.10 100.00 95.23
+Village 780 560 71.79 96.20 86.82
+TABLE V: Comparisons with other LLM-based methods, evaluated on easy subset of MultiTerraPath. $N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$.
+Method Terrain description $N_{\text{Improved}}$ $N_{\text{Equal}}$ $N_{\text{Deteriorated}}$ IR (%) $\uparrow$ SR(%) $\uparrow$ SCP(%) $\uparrow$ FPS $\uparrow$
+A* [4] (Baseline) - - - - - 100.00 86.02 9.39
+GPT-4o 208 0 792 20.80 43.10 31.23 0.08
+GPT-4-turbo 233 0 767 23.30 46.20 34.00 0.09
+Gemini-2.5-flash 48 0 952 4.80 27.10 20.26 0.03
+Claude-Opus-4 207 0 793 20.70 41.00 29.43 0.08
+LLM-A* [16] 152 538 248 69.00 93.80 87.29 0.81
+LLM-Advisor (GPT-4o) + A* 626 135 239 76.10 99.10 94.10 1.09
+LLM-Advisor (GPT-4-turbo) + A* 268 495 237 76.30 99.20 95.23 1.26
+TABLE VI: Ablation studies when using easy subset of MultiTerraPath.
+$N_{\text{Images}}=1,000$ and $N_{\text{Path}}=1,000$.
+As more structured context and grounding are introduced, LLM-Advisor produces increasingly precise and reliable suggestions, while hallucination rates consistently decrease.
+Method Terrain description Path description RAG Descriptive output $N_{\text{Suggested}}$ $N_{\text{Improved}}$ RP (%) $\uparrow$ IR (%) $\uparrow$ HR (%) $\downarrow$
 Detailed Brief
-LLM-Advisor + A* [13] - 1,000 1,000 612 432 70.59
-LLM-Advisor + RRT* [22] - 1,000 1,000 737 512 69.47
-LLM-Advisor + LLM-A* [34] - 1,000 1,000 770 606 78.70
-TABLE IV: Evaluation of planned paths across different methods on the RUGD_v2 dataset using the proposed LLM-Advisor
-Method Scenes $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
-LLM-Advisor + A* Creek 1000 1000 876 580 66.21
-Park 1000 1000 860 720 83.72
-Trail 1000 1000 620 540 87.10
-Village 1000 1000 780 560 71.79
-TABLE V: Comparison of our LLM-Advisor with other LLM-based methods
-Method Terrain description $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Successful}}$ $N_{\text{Improved}}$ $N_{\text{Equal}}$ $N_{\text{Deteriorated}}$ $IR$ (%) $\uparrow$ FPS $\uparrow$
-A* [13] - 1,000 1,000 1,000 - - - - 9.39
-GPT-4o 1,000 1,000 1,000 208 0 792 20.80 1.21
-LLM-A* [34] 1,000 1,000 938 152 538 248 69.00 3.10
-LLM-Advisor + A* 1,000 1,000 1,000 432 393 175 82.50 5.86
-Figure 9: Visualization of the planned paths by GPT-4o, A*, LLM-A*, and LLM-Advisor + A*.
-Black bold numbers indicate the favorable suggestions, while red bold numbers highlight less favorable suggestions.
-All other numbers represent cases where no suggestions were made.
-Lower costs indicate better performance
-Figure 10: Visualization of experimental results on RUGD_v2 dataset
-TABLE VI: Ablation studies
-Method Terrain description Path description RAG Descriptive output $N_{\text{Images}}$ $N_{\text{Path}}$ $N_{\text{Suggested}}$ $N_{\text{Improved}}$ $RP$ (%) $\uparrow$
-Detailed Brief
-LLM-Advisor + A* - - - 1,000 1,000 602 130 21.59
-- - - 1,000 1,000 607 352 57.99
-- - 1,000 1,000 610 413 67.70
-- - 1,000 1,000 609 399 65.52
-- 1,000 1,000 612 432 70.59
+LLM-Advisor (GPT-4o)+ A* - - - 602 130 21.59 52.80 3.30
+- - - 607 352 57.99 74.50 2.00
+- - 610 413 67.70 80.30 1.10
+- - 609 399 65.52 79.00 1.40
+- 865 626 72.37 76.10 0.70
+Figure 8: Results on MultiTerraPath dataset with LLM-Advisor (GPT-4o) + A*.
+Figure 9: Results on RUGD_v2 dataset with LLM-Advisor (GPT-4o) + A*.
+Figure 10: Results in challenging environments: (a) detours around trees in a blurry image; (b) robustness to lighting and overexposure.
 Path description (Brief): From (88, 172) to (305, 370) the terrain cost is 1.
@@
 From (337, 414) to (337, 443) the terrain cost is 5.
-The total cost of the path is -9223372036854775399.
-Figure 11: Path description (Brief)
-Figure 12: Visualization of experimental results after integrating the proposed hallucination-mitigation strategies
-Overview of RRT*.
-RRT* is a sampling-based path planning algorithm that extends RRT [21].
-It randomly samples points and connects each new point to the nearest neighbor in the tree.
-Unlike RRT, RRT* evaluates the cost from all nearby nodes and selects the lowest-cost connection.
-After connecting, it rewires the tree whenever a more efficient path is found.
-Over time, this iterative rewiring process improves the overall path quality.
-Although time-consuming, RRT* generates reasonable paths, which is why we use our LLM-Advisor to evaluate its results, aiming to assess its effectiveness as a validation tool for precomputed reasonable paths.
-Overview of LLM-A*.
-LLM-A* leverages LLMs to provide a candidate list for the A* algorithm.
-The main idea of LLM-A* is to improve search efficiency in obstacle-avoidance path planning.
-By restricting the search to the candidate list provided by LLM, LLM-A* can significantly decrease the search time.
-The prompt used for experiments using LLM-A* [34] is given in Fig.
-8.
-Since LLM-Advisor functions as an advisor, we assess its performance using two metrics: Relative Precision ($RP$) and Improvement Ratio ($IR$).
-The definition of $RP$ is provided in Eq.
-1, and the definition of $IR$ is given in Eq.
-2.
-$RP$ represents the proportion of improved paths within the suggestions made by LLM-Advisor, while $IR$ indicates the percentage of paths that are equal to or better than the original ones after incorporating LLM-Advisor.
-$IR$ can reveal how many paths are valid and effective.
+The total cost of the path is 9223372036854775399.
+Figure 11: Path description (Brief).
+Figure 12: Results on the hard subset of MultiTerraPath.
+We visualize the original A* paths and the improved paths refined by LLM-Advisor (GPT-4o).
+LLM-Advisor shows stronger effectiveness in hard scenarios with complex terrains and obstacles.
+Figure 13: Results after applying hallucination-mitigation strategies.
+We assess the performance of LLM-Advisor using five metrics: Relative Precision (RP), Improvement Ratio (IR), Success Rate (SR), newly proposed cost-aware success weighted by path length (SCP) and hallucination rate (HR).
+1) RP represents the proportion of improved paths among all suggestions made by LLM-Advisor. The definition of RP is provided in Eq. 1.
 $RP=\frac{N_{\text{Improved}}}{N_{\text{Suggested}}},$ (1)
 where $N_{\text{Improved}}$ denotes the number of improved paths and $N_{\text{Suggested}}$ denotes the number of suggested paths.
+2) IR indicates the percentage of paths that are equal to or better than the original A*-planned paths after incorporating LLM-Advisor.
+Therefore, IR reflects how many suggested paths are valid and effective.
+The definition of IR is given in Eq.
+2.
 $IR=\frac{N_{\text{Improved}}+N_{\text{Equal}}}{N_{\text{Path}}},$ (2)
 where $N_{\text{Equal}}$ denotes the number of paths that are identical to the A*-planned path, and $N_{\text{Path}}$ represents the total number of planned paths.
+3) SR measures the proportion of planned paths that successfully reach the target without collision. It is defined as
+$SR=\frac{N_{\text{Successful}}}{N_{\text{Path}}},$ (3)
+where $N_{\text{Successful}}$ denotes the number of collision-free paths that reach the goal.
+4) SCP evaluates navigation performance based on traversal cost rather than geometric path length.
+The commonly used Success weighted by Path Length (SPL) metric assumes uniform traversal cost and evaluates path optimality purely based on path length, which becomes inadequate in environments with heterogeneous terrain costs.
+In such cases, a shorter path may incur higher traversal cost due to unfavorable terrain.
+To address this limitation, we propose SCP, defined as
+$\mathrm{SCP}=\frac{1}{N_{\text{Path}}}\sum_{i=1}^{N_{\text{Path}}}S_{i}\cdot\frac{C_{i}^{\ast}}{\max\left(C_{i},\,C_{i}^{\ast}\right)},$ (4)
+where $S_{i}\in\{0,1\}$ is a per-path success indicator, $C_{i}$ denotes the total traversal cost of the path generated by the model, and $C_{i}^{\ast}$ represents the minimum-cost path obtained by a cost-aware shortest-path planner.
+The traversal cost $C_{i}$ is computed as
+$C_{i}=\sum_{k=1}^{K_{i}}d_{k}\cdot w(t_{k}),$ (5)
+where $d_{k}$ is the length of the $k$-th path segment, $t_{k}$ denotes the terrain type traversed by the segment, and $w(\cdot)$ is the corresponding terrain-dependent cost weight.
+By explicitly modeling terrain-dependent traversal cost, SCP generalizes SPL to cost-aware navigation settings and provides a more faithful measure of navigation efficiency in heterogeneous environments.
+5) HR measures the proportion of planned paths that are physically invalid or hallucinated. It is defined as
+$HR=\frac{N_{\text{Hallucinated}}}{N_{\text{Path}}},$ (6)
+where $N_{\text{Hallucinated}}$ denotes the number of paths that violate physical or environmental constraints, such as trajectories that leave the valid navigation space (e.g., paths going into the air).
 III-B Main Results
@@
 Table III presents the main comparison results for the MultiTerraPath dataset.
-Most suggestions made by the proposed LLM-Advisor are more cost-efficient.
-It achieves an $RP$ of 70.59% when making suggestions for the A*-planned path, 69.47% for the RRT*-planned path, and 78.70% for the LLM-A*-planned path.
+Most suggestions made by the proposed LLM-Advisor (GPT-4o) are more cost-efficient.
+When evaluated using RP, SR, and SCP, LLM-Advisor (GPT-4o) demonstrates superior effectiveness in harder scenarios.
 This demonstrates that, with detailed path descriptions, LLM-Advisor can effectively serve as a reliable validation tool for planned paths generated by other methods.
-Moreover, the proposed LLM-Advisor is versatile.
+Moreover, LLM-Advisor is versatile.
 It can improve the previous approaches regardless of whether they are search-based, sampling-based, or LLM-based.
-Table IV shows experimental results for the the RUGD_v2 dataset.
-Table V presents a comparison of LLM-Advisor+A* with A*, LLM-A* [34] and GPT-4o.
-The best performance for each evaluation metric is bolded, while the second-best is underlined.
-A* successfully generated valid paths for all cases.
-Since the proposed LLM-Advisor functions as an advisor, evaluating the paths generated by A*, it also successfully validated all paths.
-GPT-4o successfully planned all paths, although most were suboptimal or invalid.
-However, it was still able to generate a path.
+Table IV shows experimental results for the RUGD_v2 dataset.
+Table VI presents a comparison of LLM-Advisor+A* with A*, GPT-4o, GPT-4-turbo, Gemini-2.5-flash, Claude-Opus-4, and LLM-A* [16].
+The best performance for each evaluation metric is bolded.
 In LLM-A [16], GPT-4o suggests a target list, providing a sequence of intermediate target states for A* to consider during path planning.
 If the generated target list contains points within the obstacle region, the neighbor search process cannot proceed.
-As shown in Table V, there were 62 paths that were not successfully planned using LLM-A*.
-When considering the Improvement Ratio (IR), LLM-Advisor+A* achieved a higher IR, indicating better performance.
-The GPT-4o has the lowest IR, indicating that GPT-4o does not have ability to do zero-shot path planning.
-Moreover, since our LLM-Advisor acts as a postprocessor, it is possible to involve humans in the verification process, allowing us to decide whether to accept the paths suggested by the LLM-Advisor.
+As shown in Table VI, there were 62 paths that were not successfully planned using LLM-A*, achieving a SR of 93.80%.
+LLM-Advisor (GPT-4o) + A* yields 626 improved paths and achieves the overall Improvement Ratio (IR) of 76.10% and a SR of 99.10%.
+LLM-Advisor (GPT-4-turbo) + A* achieves the highest IR.
+The GPT-4o, GPT-4-turbo, Gemini-2.5-flash, and Claude-Opus-4 have relatively lower IR, indicating that existing LLMs does not have ability to do zero-shot path planning.
+Moreover, since our LLM-Advisor acts as a postprocessor, it is possible to involve humans in the verification process, allowing us to decide whether to accept the paths suggested by LLM-Advisor.
 Incorporating human judgment in this process can further improve the quality of the planned paths.
-To evaluate the computational efficiency of different path planning methods, we report frames per second (FPS) for each approach in Table V.
-FPS quantifies the number of paths processed per second, where a higher value indicates greater efficiency.
-All methods are executed entirely on the CPU.
-Among them, LLM-Advisor + A* selectively makes suggestions only when an improvement is needed, minimizing unnecessary LLM computations and achieving a higher FPS compared to other LLM-based methods.
+To evaluate computational efficiency, we report frames per second (FPS) for each method in Table VI.
+All methods run entirely on the CPU.
+LLM-Advisor + A* achieves a higher FPS than other LLM-based approaches.
+In addition to runtime efficiency, we also report token efficiency for LLM-Advisor.
+Token usage is measured via the OpenAI API statistics (prompt and completion tokens).
+In our implementation, each inference call corresponds to one start–goal decision and consumes approximately 3.2k tokens, with a negligible completion length when the A* path is already cost-effective.
 Visualizations.
+As shown in Figs.
+8 and 10, our LLM-Advisor successfully suggests more cost-efficient paths.
+Figs.
+10 and 13 present experimental results in more challenging environments.
+Our model demonstrates robustness to visual degradations such as blurry images and overexposure (Fig.
+10), and is able to plan significantly more cost-effective paths in harder scenarios (Fig.
+13).
+Ablation Studies.
+We experimentally verified the difference between detailed and brief path representations (Fig.
+11).
+As shown in Table VI, the granularity in path descriptions significantly impacts final performance.
+By using detailed path descriptions, the LLM-Advisor can more accurately evaluate the A*-planned path and make a higher-quality suggestion.
+We also found that the level of detail has only a limited influence on the number of suggestions.
+This indicates that the ability of LLM-Advisor to suggest improvements to planned paths is not affected much by the path descriptions.
+Moreover, using the proposed hallucination-mitigation strategies, the LLM-Advisor achieves higher RP and lower HR.
 Fig.
-9 presents some path planning results.
-As shown in the figure, our LLM-Advisor successfully suggests more cost-efficient paths.
-Fig.
-10 provides visualization results for the RUGD_v2 dataset.
-Ablation Studies.
-Table VI shows some ablation study results.
-We evaluate LLM-Advisor with both brief and detailed path descriptions to demonstrate the effectiveness of the proposed prompt.
-One example of brief path description is given in Fig.
-11.
-As shown, the level of detail in path descriptions significantly impacts final performance.
-By using detailed path descriptions, the LLM-Advisor can more accurately evaluate the A*-planned path and make a higher-quality suggestion.
-We also found that it has only a limited influence on the number of suggestions.
-This indicates that the ability of LLM-Advisor to suggest improvements to planned paths is not affected much by the path descriptions.
-Moreover, using the proposed hallucination-mitigation strategies, the LLM-Advisor achieves higher RP.
-Fig.
-12 shows the visualization results.
+13 shows the corresponding visualization results after applying the hallucination-mitigation strategies.
 IV Future Work
-The proposed LLM-Advisor operates as a zero-shot model, relying heavily on the performance of general-purpose pretrained LLMs.
-Without task-specific training, it leverages the inherent knowledge and reasoning capabilities of these models to provide guidance and insights.
-However, its performance may be limited when dealing with domain-specific challenges.
-As a next step, we plan to fine-tune the model using task-relevant data to enhance its accuracy, adaptability, and effectiveness in path planning and related tasks.
+The proposed LLM-Advisor functions as a zero-shot model, relying heavily on the performance of general-purpose pretrained LLMs.
+Therefore, its performance may be suboptimal when dealing with domain-specific challenges.
+As a next step, we plan to fine-tune the model using task-specific data to enhance its accuracy, adaptability, and effectiveness in path planning and related applications.
+We also aim to extend its capabilities to better handle ambiguous inputs, diverse environments, and noisy terrain descriptions.
 V Conclusions
-This paper introduced a novel prompt-based approach, LLM-Advisor, which achieves high relative precision and improvement ratio in evaluating paths planned by existing methods.
-The LLM-Advisor is versatile and can be used to verify paths generated by any planing method.
-It provides suggestions when existing methods fail to identify the most cost-efficient path.
-Furthermore, we propose two strategies to mitigate hallucinations.
-Our experiments on created MultiTerraPath and RUGD_v2 datasets demonstrate the effectiveness of LLM-Advisor.
-We also experimentally found that GPT-4o struggles with effective path planning based solely on terrain descriptions.
-We believe that our work offers valuable insights for the path planning community, particularly for researchers investigating the integration of LLMs into path planning.
-Acknowledgments
-This paper is partially financially supported by JSPS KAKENHI Grant Number 24K20787, Fund to Support Female Researchers with Their Research Skills Training of The University of Tokyo, and Institute for AI and Beyond of The University of Tokyo.
+This paper introduces LLM-Advisor, a novel planner-agnostic mechanism for cost refinement that operates without modifying the underlying planner.
+We further propose two effective strategies to mitigate LLM hallucinations, and introduce the MultiTerraPath and RUGD_v2 datasets for systematic evaluation.
+Extensive experimental validation demonstrates that: (1) LLM-Advisor is versatile and can be applied to paths generated by a wide range of planning methods, achieving high relative precision and improvement ratio, and providing effective suggestions when existing planners fail to identify the most cost-efficient paths; (2) the proposed hallucination-mitigation strategies significantly reduce hallucination rates; (3) unlike end-to-end planning approaches, LLM-Advisor limits the impact of hallucinations by employing LLMs in a non-decisive advisory role; and (4) state-of-the-art LLMs, including GPT-4o, GPT-4-turbo, Gemini-2.5-Flash, and Claude-Opus-4, struggle to perform effective path planning based solely on terrain descriptions.
+We believe that this work provides valuable insights for the path planning community, particularly for researchers exploring the integration of LLMs into classical planning frameworks.